The height h (in feet) of an object that is dropped from the height of s feet is given by the formula h = s - 16t 2 , where t is the time the object has been falling. A 6 foot tall woman on a sidewalk looks directly overhead and sees a window washer drop a bottle from the 5 story. How long does she have to get out of the way? Round to the nearest tenth. (A story is 12 feet.)
Answer by ankor@dixie-net.com(22740)

The height h (in feet) of an object that is dropped from the height of s feet is given by the formula h = s - 16t 2 , where t is the time the object has been falling. A 6 foot tall man on a sidewalk looks directly overhead and sees a window washer drop a bottle from the 5 story. How long does he have to get out of the way? Round to the nearest tenth. (A story is 12 feet.)
:
The distance the bottle falls to strike the man:
(5*12) - 6 = 54 ft
s = 54
:
-16t^2 + 54 = 0
-16t^2 = -54
+16t^2 = + 54
t^2 = 54/16
t^2 = 3.375
t = sqrt(3.375)
t = 1.8 sec